1.The average rate of effusion for CO2 molecules at room temperature is 409L/s. What is the molar mass of a gas whose molecules have an average velocity of 322L/s under the same conditions?

2. A sample of helium, He effuses through a tiny pore in a balloon 6.50 times faster than does an unknown gas X. What is the molar mass of the gas?

Please show all the work for me !!!

Thank you:)
Answer #1

1.The average rate of effusion for CO2 molecules at room temperature is 409 L/s. What is the molar mass of a gas whose molecules have an average velocity of 322 L/s under the same conditions?

According to Graham,s law,

R2 / R1 = sqrt (M1 / M2)

or

322 / 409 = sqrt (44 / M1 )

or

M1 = 71.0  g / mole.

molar mass of gas = 71.0  g / mole.

2. A sample of helium, He effuses through a tiny pore in a balloon 6.50 times faster than does an unknown gas X. What is the molar mass of the gas?

R2 / R1 = sqrt (M1 / M2)

or

6.50 / 1 = sqrt (M1 / 4.00)

or

M1 = 169 g / mole.

thus

molar mass of gas = 169 g / mole.